OpenStudy (lgbasallote):

how do i get the LCM for this?? \[\frac{2x^2 - x - 1}{x^2 - 9} \times \frac{x+3}{2x + 1}\]

OpenStudy (diyadiya):

just factorise

OpenStudy (lgbasallote):

oh...any hints?

Parth (parthkohli):

You don't have to! In multiplication, you just multiply, no LCM required, or you could factorize and also I see that they can be cancelled!

OpenStudy (diyadiya):

yeah x^-9= x^2-3^2

OpenStudy (lgbasallote):

then when do you LCM??

Parth (parthkohli):

@lgbasallote You take the LCM in addition or subtraction.

OpenStudy (lgbasallote):

oh i see..sorry

OpenStudy (diyadiya):

@lgbasallote 2x^2-x-1 = ?

OpenStudy (lgbasallote):

uhmmm wait...i know this...you break down the middle term right?

OpenStudy (diyadiya):

Yes

OpenStudy (lgbasallote):

i guess -2x + x could work?

Parth (parthkohli):

Yes lgba, quadratic factoring.

Parth (parthkohli):

Yep.

OpenStudy (lgbasallote):

so 2x^2 - 2x + x + 1?

Parth (parthkohli):

Yes, now factorize by grouping.

OpenStudy (diyadiya):

right

OpenStudy (lgbasallote):

uhmmm 2(x^2 - x) + x + 1? this doesnt make sense o.O

OpenStudy (diyadiya):

take 2x out not just 2

OpenStudy (lgbasallote):

ohhh finklesauce

OpenStudy (lgbasallote):

2x(x+1) + x+ 1...that looks better

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 2x(x - 1) }\)

OpenStudy (lgbasallote):

i factor out x+1 this time right?

Parth (parthkohli):

Do you know how to factor by grouping?

OpenStudy (lgbasallote):

uhmm...idk

OpenStudy (diyadiya):

its 2x^2 - 2x + x - 1

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 2x(x - 1) - (-x - 1) }\)

OpenStudy (lgbasallote):

that still doesnt make sense o.O

OpenStudy (lgbasallote):

i cant factor out x - 1....

OpenStudy (lgbasallote):

@ParthKohli

Parth (parthkohli):

Not sure about that :/

OpenStudy (diyadiya):

2x^2 - 2x + x - 1 2x(x-1) + 1(x-1)

OpenStudy (lgbasallote):

ohhhh now that makes sense...

Parth (parthkohli):

Oh yeah I forgot that there's a positive x :P

OpenStudy (lgbasallote):

so x - 1 gets factored out.... (x-1)(2x -1)?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow (2x + 1)(x - 1) }\)

OpenStudy (lgbasallote):

oh plus

OpenStudy (diyadiya):

Yeah!! plus :D

OpenStudy (lgbasallote):

so what do i do with that now =_=

Parth (parthkohli):

I used to know a genius lgba ://

Parth (parthkohli):

Cancel out with the denominator of the other fraction.

Parth (parthkohli):

Put the factored terms in the fractions and see what you can cancel!

OpenStudy (lgbasallote):

so \[\frac{x-1}{x^2 - 9} \times x +3\]?? is that right o.O

Parth (parthkohli):

Yes!

OpenStudy (lgbasallote):

oh goodie

Parth (parthkohli):

You can also factorize x^2 - 9

OpenStudy (lgbasallote):

how??

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x^2 - 9 = x^2 - 3^2 = (x + 3)(x - 3) }\)

OpenStudy (lgbasallote):

oh

OpenStudy (lgbasallote):

x+ 3 gets canceled?

Parth (parthkohli):

Now, you can cancel x + 3 too

Parth (parthkohli):

Yes

OpenStudy (lgbasallote):

hmm..im getting \[\frac{x-1}[x+3}\]

OpenStudy (lgbasallote):

\[\frac{x-1}{x+3}\]

OpenStudy (lgbasallote):

is that right?

Parth (parthkohli):

How did you get that?

OpenStudy (lgbasallote):

i dont know :/

OpenStudy (lgbasallote):

you said to factor and cancel =_=

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow {a \over b} * {c \over d} = {ac \over cd} }\) Uh huh?

OpenStudy (lgbasallote):

???

OpenStudy (lgbasallote):

ohhh oh yeah *facepalm*

OpenStudy (diyadiya):

\[\frac{\cancel{(2x+1)}(x-1)}{\cancel{(x+3)}(x-3)} \times \frac{\cancel{(x+3)}}{\cancel{(2x+1)}}\]

Parth (parthkohli):

So., you were right ^-^